OpenStudy (anonymous):

Solve the formula 3a – 4b = c for b.

OpenStudy (solomonzelman):

\[3a-4b=c\]\[3a-4b\color{red} {-3a}=c\color{red} {-3a}\]\[-4b=c-3a\]what would be your last step?

OpenStudy (anonymous):

i have no clue i never done these:(

OpenStudy (solomonzelman):

ok, it asks you to solve for b, or to set the equation the following way\[b=....\]whatever it would be equal to. You have to eliminate b with it's coefficient from other variables and then divide both sides by the coefficient of b you will get (again)\[b=....\]

OpenStudy (solomonzelman):

See what I did till now?

OpenStudy (anonymous):

im still lost i never did these kind of problems

OpenStudy (solomonzelman):

Alright, lets do different examples then. Solve for d \[2x+d=c\] eliminate d from other variables. \[2x+d\color{red} {-2x}=c\color{red} {-2x}\]preform this subtraction.\[d=c-2x\]see?